# Weight Space Representation Learning on Diverse NeRF Architectures

### v3

@@ Abstract
Neural Radiance Fields (NeRFs) have emerged as a groundbreaking paradigm for representing 3D objects and scenes by encoding shape and appearance information into the weights of a neural network. Recent studies have demonstrated that these weights can be used as input for frameworks designed to address deep learning tasks; however, such frameworks require NeRFs to adhere to a specific, predefined architecture. In this paper, we introduce the first framework capable of processing NeRFs with diverse architectures and performing inference on architectures unseen at training time. We achieve this by training a Graph Meta-Network within an unsupervised representation learning framework, and show that a contrastive objective is conducive to obtaining an architecture-agnostic latent space. In experiments conducted across 13 NeRF architectures belonging to three families (MLPs, tri-planes, and, for the first time, hash tables), our approach demonstrates robust performance in classification, retrieval, and language tasks involving multiple architectures, even unseen at training time, while also matching or exceeding the results of existing frameworks limited to single architectures.
@@ 1 Introduction
  • To achieve these goals, we created and will publicly release several new datasets of NeRFs parameterized by diverse architectures.
@@ 4.1 Latent Space Analysis
To study the impact of different learning objectives on the organization of the resulting NeRF latent space, we apply t-SNE dimensionality reduction (Van der Maaten & Hinton, 2008) to the embeddings computed by our framework on NeRFs belonging to the test set of the ALL dataset. Fig.˜3 shows the resulting bi-dimensional plots. Some interesting patterns can be noted. When trained to minimize $\mathcal{L}_{\text{R}}$ alone (Fig.˜3, left), the encoder creates an embedding space that is clustered by class, even though class labels were not used at training time. This behavior is a byproduct of the loss, which enforces NeRFs encoding similar shapes and colors to lie nearby in the latent space. Another outcome one may expect is that the same object represented by different NeRF architectures is projected into nearby embeddings; however, this turns out not to be the case. Instead, three distinct clusters emerge for each class, each corresponding to a NeRF architecture, and these clusters are further away from each other than clusters sharing the same underlying architecture. This outcome shows that $\mathcal{L}_{\text{R}}$ alone does not directly encourage the model to align NeRF embeddings regardless of the input architecture. Conversely, when trained to minimize $\mathcal{L}_{\text{C}}$ alone (Fig.˜3, right), the encoder significantly reduces the distance between different architectures representing the same class, but also provides a considerably less distinguishable clusterization by class compared to $\mathcal{L}_{\text{R}}$ case (especially for some classes, like chairs, sofas, and tables). Finally, $\mathcal{L}_{\text{R}+\text{C}}$ (Fig.˜3, middle) strikes a balance between these latent space properties. Compared to $\mathcal{L}_{\text{R}}$, one macro-cluster per class is present, although hash tables tend to form a separate sub-cluster (which is especially evident for airplanes and watercrafts). Compared to $\mathcal{L}_{\text{C}}$, classes are more separated (although not as much as with $\mathcal{L}_{\text{R}}$), but the sub-clusterization of hash tables is more apparent. Thus, we expect $\mathcal{L}_{\text{R}}$ to be the best choice for tasks where the separation between classes is highly relevant; $\mathcal{L}_{\text{R}+\text{C}}$ and $\mathcal{L}_{\text{C}}$, on the other hand, are likely to be the most effective for tasks where strong invariance to the NeRF architecture is required, with $\mathcal{L}_{\text{R}+\text{C}}$ being preferable, as it preserves more separation between classes. For these reasons, the remainder of this section shows experimental results obtained with $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$; results for $\mathcal{L}_{\text{C}}$ are reported in App.˜C.
@@ 4.4 NeRF Captioning and Q&A
Amaduzzi et al. (2024; 2025) have previously shown that language tasks on NeRFs benefit from the holistic understanding of 3D objects captured by NeRF weights, leading their weight-processing framework (dubbed LLaNA) to outperform methods ingesting rendered views as input in NeRF captioning and NeRF Q&A tasks. To achieve these results, LLaNA uses the embeddings produced by the trained nf2vec encoder to condition an LLM belonging to the Llama 2 family (Touvron et al., 2023). To assess the effectiveness of our learned embeddings in downstream language tasks, we replaced the nf2vec encoder in LLaNA’s pipeline with our GMN encoder, trained the resulting framework on textual descriptions from the ShapeNeRF–Text dataset (Amaduzzi et al., 2024), and evaluated it on brief captioning, detailed captioning, and single-round Q&A. In the multi-architecture setting, our framework is trained on ALL (as usual), and the ALL embeddings produced by our encoder are used to train the LLM. Since the nf2vec encoder can process only MLP-based NeRFs, the original LLaNA framework cannot be applied in the multi-architecture setting. In the single-architecture setting, instead, our framework is trained on MLP, and the MLP embeddings produced by our encoder are used to train the LLM. In this setting, our results are compared with the original LLaNA framework, which is also trained on MLP. Multi-architecture results are shown in Tabs.˜13, 13 and 13, whereas single-architecture results are shown in Tabs.˜13, 13 and 13. As in Amaduzzi et al. (2024; 2025), the reported metrics are a measure of the similarity between the global embeddings of the generated and ground-truth texts, where such embeddings are produced by pre-trained Sentence-BERT (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021) encoders, respectively. Extended results including traditional n-gram-based metrics are shown in App.˜F. Two main conclusions can be drawn from Tabs.˜13, 13, 13, 13, 13 and 13: in the multi-architecture setting, our method is remarkably robust to the choice of test-time NeRF architecture; when instead restricted to MLP-based NeRFs only, our method achieves results comparable to those of the original LLaNA framework. These results further confirm the architecture-agnostic nature of our embeddings and their effectiveness in downstream tasks, even as complex as those involving multimodal NeRF–language understanding.
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 68.3 70.2
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 74.1 76.0
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 81.0 81.6
@@ Appendix B Additional Experimental Details
Training. Our encoder-decoder framework is trained end-to-end for 250 epochs with batch size 8, AdamW optimizer (Loshchilov & Hutter, 2019), one-cycle learning rate scheduler (Smith & Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$ and weight decay $1\text{\times}{10}^{-2}$. These training hyperparameters are the same for $\mathcal{L}_{\text{R}}$, $\mathcal{L}_{\text{R}+\text{C}}$, and $\mathcal{L}_{\text{C}}$. In Eq.˜1, foreground and background pixels are weighted by 0.8 and 0.2, respectively, as done in Zama Ramirez et al. (2024). In Eq.˜3, $t$ and $b$ are initialized to $10$ and 10-10, respectively, as done in Zhai et al. (2023). In Eq.˜4, $\lambda$ is set to $2\text{\times}{10}^{-2}$, as we experimentally observed that this choice leads to $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{C}}$ values with the same order of magnitude.
Classification. In our classification experiments of Sec.˜4.2, the classifier is a concatenation of 3 (linear $\rightarrow$ batch-norm $\rightarrow$ ReLU $\rightarrow$ dropout) blocks, where the hidden dimensions of the linear layers are 1024, 512, and 256, and a final linear layer at the end computes the class logits. The classifier is trained via cross-entropy loss for 150 epochs with batch size 256, AdamW optimizer (Loshchilov & Hutter, 2019), one-cycle learning rate scheduler (Smith & Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$, and weight decay $1\text{\times}{10}^{-2}$. These same network and training hyperparameters are used in the classification experiments of Zama Ramirez et al. (2024).
@@ Appendix D Implementation and Hardware
Our framework implementation is built upon the codebases by Zama Ramirez et al. (2024) (for decoder and training) and Lim et al. (2024) (for the GMN architecture and graph conversion). NeRFs belonging to HASH, $\texttt{TRI}^{\texttt{OB}}$, $\texttt{HASH}^{\texttt{OB}}$, and the unseen architectures were trained with the NerfAcc framework (Li et al., 2022), as done in Zama Ramirez et al. (2024) and Cardace et al. (2024) for MLP and TRI. We plan to publicly release our code and datasets upon paper acceptance. All our experiments were performed on a single NVIDIA RTX A6000. Training either $\mathcal{L}_{\text{R}}$ or $\mathcal{L}_{\text{R}+\text{C}}$ took ${\sim}2$ weeks, whereas training $\mathcal{L}_{\text{C}}$ took ${\sim}4$ days.
@@ Appendix F Additional NeRF Captioning and Q&A Results
Tabs.˜24, 24, 24, 24, 24 and 24 show extended results for the language tasks described in Sec.˜4.4 by adding BLEU-1 (Papineni et al., 2002), ROUGE-L (Chin-Yew, 2004), and METEOR (Banerjee & Lavie, 2005) metrics, which are traditional handcrafted similarity measures based on n-gram statistics. As discussed in Sec.˜4.4, S-BERT (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021) are instead based on pre-trained networks, and generally considered more effective than n-gram-based metrics at measuring the quality of the generated text (Amaduzzi et al., 2024).
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 68.3 70.2 33.1 35.3 36.6
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 74.1 76.0 20.1 32.5 20.6
@@
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 81.0 81.6 45.4 52.5 49.4

### v5

@@ Abstract
Neural Radiance Fields (NeRFs) have emerged as a groundbreaking paradigm for representing 3D objects and scenes by encoding shape and appearance information into the weights of a neural network. Recent studies have demonstrated that these weights can be used as input for frameworks designed to address deep learning tasks; however, such frameworks require NeRFs to adhere to a specific, predefined architecture. In this paper, we introduce the first framework capable of processing NeRFs with diverse architectures and performing inference on architectures unseen at training time. We achieve this by training a Graph Meta-Network within an unsupervised representation learning framework, and show that a contrastive objective is conducive to obtaining an architecture-agnostic latent space. In experiments conducted across 13 NeRF architectures belonging to three families (MLPs, tri-planes, and, for the first time, hash tables), our approach demonstrates robust performance in classification, retrieval, and language tasks involving multiple architectures, even unseen at training time, while also matching or exceeding the results of existing frameworks limited to single architectures. Our code and data are available at this https URL.
@@ 4.1 Latent Space Analysis
To study the impact of different learning objectives on the organization of the resulting NeRF latent space, we apply t-SNE dimensionality reduction (Van der Maaten and Hinton, 2008) to the embeddings computed by our framework on NeRFs belonging to the test set of the ALL dataset. Fig.˜3 shows the resulting bi-dimensional plots. Some interesting patterns can be noted. When trained to minimize $\mathcal{L}_{\text{R}}$ alone (Fig.˜3, left), the encoder creates an embedding space that is clustered by class, even though class labels were not used at training time. This behavior is a byproduct of the loss, which enforces NeRFs encoding similar shapes and colors to lie nearby in the latent space. Another outcome one may expect is that the same object represented by different NeRF architectures is projected into nearby embeddings; however, this turns out not to be the case. Instead, three distinct clusters emerge for each class, each corresponding to a NeRF architecture, and these clusters are further away from each other than clusters sharing the same underlying architecture. This outcome shows that $\mathcal{L}_{\text{R}}$ alone does not directly encourage the model to align NeRF embeddings regardless of the input architecture. Conversely, when trained to minimize $\mathcal{L}_{\text{C}}$ alone (Fig.˜3, right), the encoder significantly reduces the distance between different architectures representing the same class, but also provides a considerably less distinguishable clusterization by class compared to $\mathcal{L}_{\text{R}}$ case (especially for some classes, like chairs, sofas, and tables). Finally, $\mathcal{L}_{\text{R}+\text{C}}$ (Fig.˜3, middle) strikes a balance between these latent space properties. Compared to $\mathcal{L}_{\text{R}}$, one macro-cluster per class is present, although hash tables tend to form a separate sub-cluster (which is especially evident for airplanes and watercrafts). Compared to $\mathcal{L}_{\text{C}}$, classes are more separated (although not as much as with $\mathcal{L}_{\text{R}}$), but the sub-clusterization of hash tables is more apparent. Thus, we expect $\mathcal{L}_{\text{R}}$ to be the best choice for tasks where the separation between classes is highly relevant; $\mathcal{L}_{\text{R}+\text{C}}$ and $\mathcal{L}_{\text{C}}$, on the other hand, are likely to be the most effective for tasks where strong invariance to the NeRF architecture is required, with $\mathcal{L}_{\text{R}+\text{C}}$ being preferable, as it preserves more separation between classes. For these reasons, the remainder of this section shows experimental results obtained with $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$; results for $\mathcal{L}_{\text{C}}$ are reported in App.˜C.
@@ 4.4 NeRF Captioning and Q&A
Amaduzzi et al. (2024; 2025) have previously shown that language tasks on NeRFs benefit from the holistic understanding of 3D objects captured by NeRF weights, leading their weight-processing framework (dubbed LLaNA) to outperform methods ingesting rendered views as input in NeRF captioning and NeRF Q&A tasks. To achieve these results, LLaNA uses the embeddings produced by the trained nf2vec encoder to condition an LLM belonging to the Llama 2 family (Touvron et al., 2023). To assess the effectiveness of our learned embeddings in downstream language tasks, we replaced the nf2vec encoder in LLaNA’s pipeline with our GMN encoder, trained the resulting framework on textual descriptions from the ShapeNeRF–Text dataset (Amaduzzi et al., 2024), and evaluated it on brief captioning, detailed captioning, and single-round Q&A. In the multi-architecture setting, our framework is trained on ALL (as usual), and the ALL embeddings produced by our encoder are used to train the LLM. Since the nf2vec encoder can process only MLP-based NeRFs, the original LLaNA framework cannot be applied in the multi-architecture setting. In the single-architecture setting, instead, our framework is trained on MLP, and the MLP embeddings produced by our encoder are used to train the LLM. In this setting, our results are compared with the original LLaNA framework, which is also trained on MLP. Multi-architecture results are shown in Tabs.˜13, 13 and 13, whereas single-architecture results are shown in Tabs.˜13, 13 and 13. As in Amaduzzi et al. (2024; 2025), the reported metrics are a measure of the similarity between the global embeddings of the generated and ground-truth texts, where such embeddings are produced by pre-trained Sentence-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) encoders, respectively. Extended results including traditional n-gram-based metrics are shown in App.˜F. Two main conclusions can be drawn from Tabs.˜13, 13, 13, 13, 13 and 13: in the multi-architecture setting, our method is remarkably robust to the choice of test-time NeRF architecture; when instead restricted to MLP-based NeRFs only, our method achieves results comparable to those of the original LLaNA framework. These results further confirm the architecture-agnostic nature of our embeddings and their effectiveness in downstream tasks, even as complex as those involving multimodal NeRF–language understanding.
@@
$\mathcal{L}_{\text{R}}$ (ours) 68.3 70.2
@@
$\mathcal{L}_{\text{R}}$ (ours) 74.1 76.0
@@
$\mathcal{L}_{\text{R}}$ (ours) 81.0 81.6
@@ Appendix B Additional Experimental Details
Training. Our encoder-decoder framework is trained end-to-end for 250 epochs with batch size 8, AdamW optimizer (Loshchilov and Hutter, 2019), one-cycle learning rate scheduler (Smith and Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$ and weight decay $1\text{\times}{10}^{-2}$. These training hyperparameters are the same for $\mathcal{L}_{\text{R}}$, $\mathcal{L}_{\text{R}+\text{C}}$, and $\mathcal{L}_{\text{C}}$. In Eq.˜1, foreground and background pixels are weighted by 0.8 and 0.2, respectively, as done in Zama Ramirez et al. (2024). In Eq.˜3, $t$ and $b$ are initialized to $10$ and 10-10, respectively, as done in Zhai et al. (2023). In Eq.˜4, $\lambda$ is set to $2\text{\times}{10}^{-2}$, as we experimentally observed that this choice leads to $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{C}}$ values with the same order of magnitude.
Classification. In our classification experiments of Sec.˜4.2, the classifier is a concatenation of 3 (linear $\rightarrow$ batch-norm $\rightarrow$ ReLU $\rightarrow$ dropout) blocks, where the hidden dimensions of the linear layers are 1024, 512, and 256, and a final linear layer at the end computes the class logits. The classifier is trained via cross-entropy loss for 150 epochs with batch size 256, AdamW optimizer (Loshchilov and Hutter, 2019), one-cycle learning rate scheduler (Smith and Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$, and weight decay $1\text{\times}{10}^{-2}$. These same network and training hyperparameters are used in the classification experiments of Zama Ramirez et al. (2024).
@@ Appendix D Implementation and Hardware
Our framework implementation is built upon the codebases by Zama Ramirez et al. (2024) (for the decoder and training) and Lim et al. (2024) (for the GMN architecture and graph conversion). NeRFs belonging to HASH, $\texttt{TRI}^{\texttt{OB}}$, $\texttt{HASH}^{\texttt{OB}}$, and the unseen architectures were trained with the NerfAcc framework (Li et al., 2022), as done in Zama Ramirez et al. (2024) and Cardace et al. (2024) for MLP and TRI. All our experiments were performed on a single NVIDIA RTX A6000. Training either $\mathcal{L}_{\text{R}}$ or $\mathcal{L}_{\text{R}+\text{C}}$ took ${\sim}2$ weeks, whereas training $\mathcal{L}_{\text{C}}$ took ${\sim}4$ days.
@@ Appendix F Additional NeRF Captioning and Q&A Results
Tabs.˜24, 24, 24, 24, 24 and 24 show extended results for the language tasks described in Sec.˜4.4 by adding BLEU-1 (Papineni et al., 2002), ROUGE-L (Chin-Yew, 2004), and METEOR (Banerjee and Lavie, 2005) metrics, which are traditional handcrafted similarity measures based on n-gram statistics. As discussed in Sec.˜4.4, S-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) are instead based on pre-trained networks, and generally considered more effective than n-gram-based metrics at measuring the quality of the generated text (Amaduzzi et al., 2024).
@@
$\mathcal{L}_{\text{R}}$ (ours) 68.3 70.2 33.1 35.3 36.6
@@
$\mathcal{L}_{\text{R}}$ (ours) 74.1 76.0 20.1 32.5 20.6
@@
$\mathcal{L}_{\text{R}}$ (ours) 81.0 81.6 45.4 52.5 49.4